Qx = 235 - 3Px + 40A - 20U + 8Pw
The number of detailing jobs he gets per month (Qx) depends on the price he charges per job (Px), his monthly advertising expenditures (A) measured in Rs. 1,000s, the regional percentage unemployment rate (U), and the average price charged by local car wash businesses (Pw) for a standard wash and wax.
c) Rahim is currently charging Rs 65 per detailing job and spending Rs 3,500 per month on advertising. The regional unemployment rate is 7.5% and the average price of a wash and wax at a local car wash is Rs 15. How many detailing jobs per month can Rahim expect under these conditions?
(e) Calculate the Advertising and price elasticity of demand under conditions described in part c. Is the demand elastic or inelastic? Also calculate the total and marginal revenue.
1
Expert's answer
2017-11-30T15:37:07-0500
Qx = 235 - 3Px + 40A - 20U + 8Pw.
(c) A = Rs 65 per detailing job, A = Rs 3,500 per month, U = 7.5%, Pw = Rs 15.
Rahim expect under these conditions:
Qx = 235 - 3*65 + 40*3,500 - 20*0.075 + 8*15 = 140,158.5 units.
(e) Price elasticity of demand is Ed = -3, so the demand is elastic.
The total and marginal revenue are:
TR = P*Q = 65*140,158.5 = Rs 9,110,302.5.
MR = P = Rs 65.